# An Extension of the Euler-Maclaurin Summation Formula to Nearly Singular Functions

### v1

@@ 1 Introduction
at a target point $\tilde{\mathbf{x}}$ that is at a distance $d\approx 0$ from the surface $S$, which leads to near-singularity of the type (1) in the kernel $K(\tilde{\mathbf{x}},\mathbf{x})$ [7].
@@
In this paper, we provide a new extension of the Euler-Maclaurin formula for nearly singular integrals of the form (1), which is derived from a combination of the above-mentioned singular Euler-Maclaurin formulas and the rational zeta series that draw connections between special functions. One unexpected consequence is an “off-mesh” Euler-Maclaurin formula for a hypersingular (i.e. finite-part) integral that is not seen in the literature before (see Corollary 10). It is also worth mentioning that the current work is closely connected to the work of Nitsche [13], in which the near-singular error corrections for the Trapezoidal rule are computed numerically (instead of using the analytic Euler-Maclaurin type expressions).
@@ 2 Preliminary: Euler-Maclaurin formulas for singular functions
where $h=\frac{a}{n}$, and where $C_{h}^{-a}$ and $C_{h}^{a}$ are edge corrections at $x=-a$ and $x=a$, respectively, which contain sufficiently many correction terms from the classical Euler-Maclaurin formula or the Gregory interpolants [5], such that the dominant quadrature error comes from the singularity of $f$ at $x=0$. The error of the punctured Trapezoidal rule is
@@
$L_{h,s}[f]$ $:=\sum_{k=-n}^{-1}f(k\,h+s\,h)\,h+C_{h}^{-a+s}$ (6)
$R_{h,s}[f]$ $:=\sum_{k=1}^{n}f(k\,h+s\,h)\,h+C_{h}^{a+s},$
$T_{h,s}[f]$ $:=L_{h,s}[f]+R_{h,s}[f].$
where the shift is $s\,h$ for $|s|\leq\frac{1}{2}$; note that when $s=0$, (6) reduces to (4).
@@
where “$f.p.$” indicates that an integral is understood as a Hadamard finite-part integral, and where $\zeta_{h}(z,s)$ is a modified Hurwitz zeta function given by
@@
where $\zeta(z,s)$ is the classical Hurwitz zeta function, and where $\psi(s)$ is the digamma function.
@@
As a special case of Theorem 1, when $s=0$ we have the following corollary, whose proof can be found in [16].
@@ 3 An Euler-Maclaurin formula for near-singular functions
where $c,d>0$ are constants, $g(x)$ a smooth function compactly supported on $[-a,a]$, and $x_{s}\in(-a,a)$ the location of near singularity. We will analyze the error of the punctured Trapezoidal rule $T_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}$; see Figure 1 for a schematic of the punctured Trapezoidal rule. Our analysis begins with the case $x_{s}=0$ such that the near-singularity is at a mesh point, namely $x=0$. Then we will consider the general case of $x_{s}\neq 0$. In both cases, we will show that the error $E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}$ consists of two components: a “singular error” which is a continuous extension of the generalized Euler-Maclaurin formulas from Section 2, and a “jump error” associated with the discontinuity of the value of the integral with respect to $d$.
Figure 1: Schematic of the punctured trapezoidal rule of mesh size $h$ for the near-singular integral (14), which can arise from an integral operator (2) with its target point $\tilde{\mathbf{x}}$ located at a distance $d$ from the surface $S$, whose orthogonal projection onto $S$ corresponds to $x_{s}=s\cdot h$ in the parameter space relative to the closest mesh point $x=0$.
@@ 3.1 Near-singular quadrature error analysis
Given $g(x)\in C^{\infty}([a,a])$ and $c,d>0$, and using the notations (3-5), we have
@@
Substituting (21,23,24) into (20) yields (17). Note that for fixed $h$, both sides of (17) are analytic functions of $\lambda=\frac{d}{ch}$, thus the identity also holds for $\lambda\geq 1$ by analytic continuation.
@@
Lemma 5.
Define the family of rational zeta series
$f_{0}(z)$ $=\tfrac{\psi(1+z)-\psi(1-z)}{2z},$ (26)
$f_{1}(z)$ $=-\tfrac{\psi(1+z)+\psi(1-z)}{2}-\log h,$
$f_{k}(z)$ $=\zeta(2-k)+z^{2}f_{k-2}(z),$ $k>1.$
Proof.
@@
This agrees with the finite-part quadrature correction of [17, eq.(3.2)-(3.4)] :
@@ 3.2 Euler-Maclaurin formula for an off-mesh near singularity
Note that $\lim_{s\to 0}p_{k,s}=z_{k}+(-1)^{k}z_{k}$, so the formula (33) reduces to (15) as $x_{s}\to 0$.
@@
Using the notations (LABEL:eq:notation_int,eq:notation_trap_shift), we have, for any integer $k\geq 0$,
@@
The connection between the rational zeta series (39) and the digamma function $\psi$ can be established by the relation [14, eq.(25.11.12)]
@@
then substituting (LABEL:eq:lemma_right_nearsingEM_shifted,eq:lemma_left_nearsingEM_shifted) into the above series yields
@@ 3.3 New singular Trapezoidal rule for off-mesh singularity
The following quadrature for a finite-part integral is presented in [17, eq.(3.10)]
@@
which is spectrally accurate; here, $\tilde{T}$ represents an ordinary Trapezoidal rule (not punctured). One can derive this formula as a consequence of Corollary 10. Let $s\to 1/2$ (hence $x_{s}\to h/2$) in (50) and using the fact that [14, eq.(25.11.8), eq.(5.15.5)]

### v2

@@ 1 Introduction
at a target point $\tilde{\mathbf{x}}$ that is at a distance $d\approx 0$ from the surface $S$, which leads to near-singularity of the type (1) in the kernel $K(\tilde{\mathbf{x}},\mathbf{x})$ [8]. The accurate and efficient evaluation of near-singular integrals is important in numerical simulations of physical systems that involve closely interacting components, such as viscous particulate flows or potential flows [3, 14].
@@
In this paper, we provide a new extension of the Euler-Maclaurin formula for nearly singular integrals of the form (1), which is derived from a combination of the above-mentioned singular Euler-Maclaurin formulas and the rational zeta series that draw connections between special functions. One unexpected consequence is an “off-mesh” Euler-Maclaurin formula for a hypersingular (i.e. finite-part) integral that is not seen in the literature before (see Corollary 10). It is also worth mentioning that the current work is closely connected to the work of Nitsche [14], in which the near-singular error corrections for the Trapezoidal rule are computed numerically (instead of using the analytic Euler-Maclaurin type expressions). While the results of the current paper provide more theoretical insights into the error analysis and makes the error correction easier to compute, the approach of Nitsche [14] can be more easily generalized to integrals in higher dimensions and has recently been applied to simulations of viscous fluid flows past ellipsoidal objects [15].
@@ 2 Preliminary: Euler-Maclaurin formulas for singular functions
where $h=\frac{a}{n}$, and where $C_{h}^{-a}$ and $C_{h}^{a}$ are edge corrections at x=ax=-a and $x=a$, respectively, from the classical Euler-Maclaurin formula. For example, $C_{h}^{a}$ has the form
$C_{h}^{a}=-\frac{1}{2}f(a)h-\sum_{k=1}^{\lfloor p/2\rfloor}\frac{B_{2k}}{(2k)!}f^{(2k-1)}(a)h^{2k}$ (5)
where $B_{2k}$ are the Bernoulli numbers and where $p$ is sufficiently large such that the dominant quadrature error comes from the singularity of $f$ at $x=0$; alternatively, $C_{h}^{a}$ can also be approximated using Gregory interpolants [6]. An implementation of the Gregory interpolants is available in the Matlab code for the numerical examples in this paper, see Remark 5.
The error of the punctured Trapezoidal rule is
@@
$L_{h,s}[f]$ $=\sum_{k=-n}^{-1}f(k\,h+s\,h)\,h+C_{h}^{-a+s}$ (7)
$R_{h,s}[f]$ $=\sum_{k=1}^{n}f(k\,h+s\,h)\,h+C_{h}^{a+s},$
$T_{h,s}[f]$ $=L_{h,s}[f]+R_{h,s}[f].$
where the shift is shs\,h for $|s|\leq\frac{1}{2}$; note that when $s=0$, (7) reduces to (4).
@@
where “f.p.f.p.” indicates that an integral is understood as a Hadamard finite-part integral, and where $\zeta_{h}(z,s)$ is a modified Hurwitz zeta function given by
@@
where $\zeta(z,s)=\sum_{n=0}^{\infty}(n+s)^{-z}$ is the classical Hurwitz zeta function (which has an analytic extension for all $z\in\mathbb{C}\setminus\{1\}$), and where $\psi(s)={\Gamma^{\prime}(s)}/{\Gamma(s)}$ is the digamma function and $\Gamma(s)$ is the Euler’s gamma function.
@@
As a special case of Theorem 1, when $s=1$ we have the following corollary, whose proof can be found in [18].
@@ 3 An Euler-Maclaurin formula for near-singular functions
where $c,d>0$ are constants, $g(x)$ a smooth function compactly supported on [a,a][-a,a], and $x_{s}\in(-a,a)$ the location of near singularity. We will analyze the error of the punctured Trapezoidal rule $T_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}$; see Figure 1 for a schematic of the punctured Trapezoidal rule. Our analysis begins with the case $x_{s}=0$ such that the near-singularity is at a mesh point, namely $x=0$. Then we will consider the general case of $x_{s}\neq 0$. In both cases, we will show that the error $E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}$ consists of two components: a “singular error” which is a continuous extension of the generalized Euler-Maclaurin formulas from Section 2, and a “jump error” associated with the discontinuity of the value of the integral with respect to $d$.
Figure 1: Schematic of the punctured trapezoidal rule of mesh size $h$ for the near-singular integral (15), which can arise from an integral operator (2) with its target point $\tilde{\mathbf{x}}$ located at a distance $d$ from the surface $S$, whose orthogonal projection onto $S$ corresponds to $x_{s}=s\cdot h$ in the parameter space relative to the closest mesh point $x=0$. The open circle at $x=0$ indicates its omission in the summation of the punctured Trapezoidal rule.
@@ 3.1 Near-singular quadrature error analysis
Given $g(x)\in C^{\infty}([a,a])$ and $c,d>0$, and using the notations (3-6), we have
@@
Lemma 4.
Define the family of rational zeta series
$f_{0}(z)$ $=\tfrac{\psi(1+z)-\psi(1-z)}{2z},$ (19)
$f_{1}(z)$ $=-\tfrac{\psi(1+z)+\psi(1-z)}{2}-\log h,$
$f_{k}(z)$ $=\zeta(2-k)+z^{2}f_{k-2}(z),$ $k>1.$
Proof.
@@
Substituting (24,26,27) into (23) yields (20). Note that for fixed $h$, both sides of (20) are analytic functions of $\lambda=\frac{d}{ch}$, thus the identity also holds for $\lambda\geq 1$ by analytic continuation.
@@
This agrees with the finite-part quadrature correction of [19, eq.(3.2)-(3.4)] :
@@ 3.2 Euler-Maclaurin formula for an off-mesh near singularity
Note that $\lim_{s\to 0}p_{k,s}=z_{k}+(-1)^{k}z_{k}$, so the formula (34) reduces to (16) as $x_{s}\to 0$. In order to prove Theorem 6, we first need the next three lemmas.
@@
Using the notations (3) and (7), we have, for any integer $k\geq 0$,
@@
The connection between the rational zeta series (40) and the digamma function $\psi$ can be established by the relation [16, eq.(25.11.12)]
@@
then substituting (3637) into the above series yields
@@ 3.3 New singular Trapezoidal rule for off-mesh singularity
The following quadrature for a finite-part integral is presented in [19, eq.(3.10)]
@@
which is spectrally accurate; here, $\tilde{T}$ represents an ordinary Trapezoidal rule (not punctured). One can derive this formula as a consequence of Corollary 10. Let $s\to 1/2$ (hence $x_{s}\to h/2$) in (51) and using the fact that [16, eq.(25.11.8), eq.(5.15.5)]
@@ 4 Numerical Examples
Remark 5.
For a general integral where the analytic expression of $g(x)$ is unknown or where $g(x)$ cannot be analytically continued to a neighborhood of the near-singular point, a truncation of the error corrections (16) and (34) can be computed by approximating $g^{(k)}(0)$ or $g^{(k)}(x_{s})$ using finite difference formulas on a small stencil centered at $x=0$. In practice, we find that convergence results that are almost identical to Figures 2 and 3 can be achieved by including correction terms in (16) and (34) that involve up to the $6^{\rm th}$ derivative of $g(x)$. Matlab code for these computations are available at https://github.com/bobbielf2/PolygammaTrap.